A 1000-g ball is tied to a string 60 cm long and is whirled in a vertical circle at 1 revolution per second. If the tension at the highest point is zero, what is the speed of the body at the highest point?


1
Expert's answer
2021-05-20T12:39:34-0400

Let's write the tension at the highest point:


"F_T=F_c-mg,""0=\dfrac{mv^2}{r}-mg,""\dfrac{mv^2}{r}=mg,""v=\sqrt{gr}=\sqrt{9.8\ \dfrac{m}{s^2}\cdot0.6\ m}=2.42\ \dfrac{m}{s^2}."